OpenStudy (anonymous):

find x from the equation x cot(90°+A)+tan(90°+A) sin A +csc(90°+A) = 0

OpenStudy (shamim):

x cot(90+A)+tan(90+A)sinA+csc(90+A)=0 -x tan A-cot A sinA+sec A=0 -x tan A-cos A+sec A=0 -x sinA/cos A-cos A+1/cos A=0 (-x sin A- cos^2 A+1)/cos A=0 -x sin A+1-cos^2A=0 -x sin A+sin ^2 A=0 sin A(-x+sin A)=0 sin A=0 A=0 or. x= sin A